(a) Find the number of relations on the set S={a, b, c, d, e}?
(b) How many relations are there on the set S={a, b, c, d, e} that contain (a, a) and (b, b)?
1
Expert's answer
2020-12-01T06:30:56-0500

We shall use the fact that the number of all subsets of "n"-element set is "2^n."


(a) Let "S=\{a,b,c,d,e\}." By defenition, a biniry relation "\rho" on a set "S" is a subset of a Cartesian square "S\times S." Since "|S\times S|=5\cdot 5=25", there are "2^{25}=33,554,432" binary relations on the set "S=\{a,b,c,d,e\}."


(b) Let us find the number of binary relations "\rho" on the set "S=\{a,b,c,d,e\}" that contain "(a, a)" and "(b, b)". Since elements of a relation "\rho\setminus\{(a,a),(b,b)\}" can be arbitrary elements of "S\times S\setminus\{(a,a),(b,b)\}", the number of binary relation on "S" that contain "(a, a)" and "(b, b)" is equal to the number of binary relations "R\subset S\times S\setminus\{(a,a),(b,b)\}". Since "|S\times S\setminus\{(a,a),(b,b)\}|=25-2=23" , there are "2^{23}=8,388,608" binary relations on the set "S=\{a,b,c,d,e\}" that contain "(a, a)" and "(b, b)".